If f(x)+f (x+a) + f (x+2a) + …. +f(x+na) = constant; ¥xεR and a> 0 and f(x) is periodic, then period of f(x) is


 

Grade:Upto college level

2 Answers

Aman Bansal
12 years ago

Dear Ranga,

Its given that 

 

 f(x)+f (x+a) + f (x+2a) + …. +f(x+na) = k(constant)

Also ,

f (x+a) + f (x+2a) + …. +f(x+na)+f(x=(n+1)a) = k(constant)

 

subtracting eq. 2 from 1:

=> F(x)= f(x+(n+1)a)

Thus period is (n+1)a

Best Of luck

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

aakanksha uday
12 years ago

hi aman

take f(x) as a periodic fnc. like sinx and a as constan quantity

it will give u ur ans...

wish u luck